OpenStudy (anonymous):

eq. of a circle: center at origin, radius is 3/4

Directrix (directrix):

Attached is the standard form for the equation of a circle.

Directrix (directrix):

Your (h,k) = (0,0) and your radius = 3/4 so post what you think the equation is. Then, we can compare answers, okay?

OpenStudy (anonymous):

x^2+y^2=9/16

Directrix (directrix):

Correct. Do you know if you are supposed to leave the equation in standard form? I was just wondering.

OpenStudy (anonymous):

nope mehehe

Directrix (directrix):

Okay.